Grade 11Mechanics

a body is moving along a straight line its distance Xt from a point on its path at time t after passing that point is given by
Xt = 8t2 -3t3 where Xt and t is in S.I. unit what is the average speed during time intarval t=0 to t=4

sachidanand
8 Years agoGrade 11

1 Answer

amit kumar patel
8 Years ago
Xt = 8t2 -3t3X(0)=0X(4)=8×14-3×64=112-192=80As we know that average velocity equal to rate of change of displacement.Vavg=X(4)-X(0)/(4-0) =80/4 =20m/s